In a hybrid corn research project, 200 seeds were planned, and 170 of them germinated. Find the empirical probability that any particular seed of this type will germinate.
My answer is P(E)= 200/170 however I don't believe I am understanding this

Answer by jim_thompson5910(35256)

P(Germinate) = # that germinated/# total = 170/200 = 17/20 = 0.85


So there's an 85% chance that a seed will germinate